OpenStudy (anonymous):

The function f(x) is shown below. f(x) = (2)x+1 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? f(x) = (2)x + 13 f(x) = (2)x + 14 f(x) = (2)x -12 f(x) = (2)x - 13

OpenStudy (anonymous):

none of them? these are shifting up and down D:

OpenStudy (anonymous):

unless it's written like 2(x+13) f(x) = 2(x+14) f(x) = 2(x-12)

OpenStudy (anonymous):

I put the wrong ones. 0.0 f(x) = (3)x + 11 f(x) = (3)x + 13 f(x) = (3)x -15 f(x) = (3)x - 13 But they're written exactly like this.

OpenStudy (anonymous):

for the second problem, you need to evaluate the problem using f(x+13) That would make it equate to 3^ (x+13-2) = f(x+13) = 3^(x+11) for the power, add the ^ sign. I mistaken in for multiplication haha.